Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 125     Entry time: 07 Apr 2023, 23:00
Author: Alejandro Laso Garcia 
Group: HED 
Subject: HED Shift Summary 07.04. 

Morning - Users from proposal 3455 continued with their measurement program. Beam stable during the measurements.

Afternoon - Changed the setup from 3455 to 3326.

Evening - Beginning of the alignment for users. Beam stable.

ELOG V3.1.4-7c3fd00